10x = 2y - 7
y=-5x + 3

HELP PLEASE!!!

Answer by stanbon(75887)

10x = 2y - 7
y=-5x + 3
---------------
Rewrite each equation:
10x - 2y = -7
-5x - y = -3
-----------------
Multiply the 2nd equation by 2:
10x - 2y = -7
-10x -2y = -6
-------------------------------------
Add the equations to get:
-4y = -13
y = 13/4
---------------
Substitute into 10x = 2y -7 to solve for "x":
10x = 2(13/4) -7
10x = 13/2 - 14/2
10x = -1/2
x = -1/20
-------------------
Cheers,
Stan H.